# NeuralSurv: Deep Survival Analysis with Bayesian Uncertainty Quantification

### v1

@@ Appendix I Computational Speed-Ups
Fix an arbitrary index $i\in\{1,\dots,N\}$. From Section 2.2, recall that $p_{\boldsymbol{\theta}}(\boldsymbol{\theta})$ is the probability density function of a multivariate normal distribution with zero mean and identity covariance matrix $\mathbf{I}_{m}$. Our goal is to show that the normalizing constant $Z(t,\mathbf{x}_{i})$ admits a strictly positive lower bound on $[0,y_{i}]$, from which the integrability of $\lambda_{0}(t,\mathbf{x}_{i};\phi)$ will follow.
Step 1: Continuity of $Z(t,\mathbf{x}_{i})$ on $[0,y_{i}]$.
Fix any $t_{0}\in[0,y_{i}]$, and let $(t_{n})_{n\geq 1}$ be a sequence in $[0,y_{i}]$ such that $t_{n}\to t_{0}$ as $n\to\infty$. Define, for each $n$, the functions
$h_{n}(\boldsymbol{\theta})$ $:=\sigma(g(t_{n},\mathbf{x}_{i};\boldsymbol{\theta}))p_{\boldsymbol{\theta}}(\boldsymbol{\theta}),\quad n\geq 1,$
$h(\boldsymbol{\theta})$ $:=\sigma(g(t_{0},\mathbf{x}_{i};\boldsymbol{\theta}))p_{\boldsymbol{\theta}}(\boldsymbol{\theta}).$
@@
since $10<\sigma(\cdot)<$ 1. Because $p_{\boldsymbol{\theta}}(\boldsymbol{\theta})$ integrates to 1 over $\mathbb{R}^{m}$, we may apply the Dominated Convergence Theorem (DCT) to conclude that:
@@
Since $t_{0}$ was arbitrary in $[0,y_{i}]$, $Z$ is continuous everywhere on that interval.
Step 2: Strict positivity of $Z(t,\mathbf{x}_{i})$ on $[0,y_{i}]$.
@@
Since $Z(t,\mathbf{x}_{i})$ is a continuous and strictly positive function on the compact interval $[0,y_{i}]$, the Weierstrass Extreme Value Theorem ensures that $Z$ attains a minimum on this interval. Define:
@@
It is straightforward to verify that $\lambda_{0}(t;\phi)$ is integrable on $[0,y_{i}]$, therefore it follows that
@@
$\mathbb{E}_{\omega\sim p_{\text{PG}}(\omega|1,0)}[\omega]=\frac{1}{4},$ (48)
which follows by taking the limit $c\to 0$ in equation (22). Alternatively, to prove (48), one can start from the representation in equation (19), apply Tonelli’s theorem to interchange expectation and infinite summation, and then invoke the series identity
@@
$\int_{0}^{y_{i}}\int_{0}^{\infty}\min(|f(\omega,-g(t,\mathbf{x}_{i};\boldsymbol{\theta}))|,1)\lambda_{i}(t,\omega;\phi)\mathrm{d}\omega\mathrm{d}t<\infty.$ (49)
@@
$\mathcal{I}_{1}$ $:=\int_{0}^{y_{i}}\int_{0}^{\infty}\frac{|g(t,\mathbf{x}_{i};\boldsymbol{\theta})|}{2}\lambda_{i}(t,\omega;\phi)\mathrm{d}t\mathrm{d}\omega,$
$\mathcal{I}_{2}$ $:=\int_{0}^{y_{i}}\int_{0}^{\infty}\frac{g(t,\mathbf{x}_{i};\boldsymbol{\theta})^{2}}{2}\omega\lambda_{i}(t,\omega;\phi)\mathrm{d}\omega\mathrm{d}t,$
$\mathcal{I}_{3}$ $:=\log(2)\int_{0}^{y_{i}}\int_{0}^{\infty}\lambda_{i}(t,\omega;\phi)\mathrm{d}\omega\mathrm{d}t.$
@@
Since $g(t,\mathbf{x}_{i};\boldsymbol{\theta})$ is continuous on the compact interval $[0,y_{i}]$, it is bounded by some $M>0$. Then,
@@
Likewise $g(t,\mathbf{x}_{i};\boldsymbol{\theta})^{2}$ is bounded by some $C>0$ over $[0,y_{i}]$ and $\mathbb{E}_{\omega\sim p_{\text{PG}}(\omega|1,0)}[\omega]=\frac{1}{4}$ (see  (48)), so
@@
Since $\mathcal{I}_{1},\mathcal{I}_{2},\mathcal{I}_{3}$, are all finite, the condition in (49) is satisfied and the sum $H(\Psi_{i})$ converges absolutely with probability 1. ∎ The next result presents an integral identity which is key to proving the data augmentation scheme of Theorem 3.1.
@@
$\int_{0}^{y_{i}}\int_{0}^{\infty}\left(1-e^{f(\omega,-g(t,\mathbf{x}_{i};\boldsymbol{\theta}))}\right)p_{\text{PG}}(\omega|1,0)\lambda_{0}(t,\mathbf{x}_{i};\phi)\mathrm{d}\omega\mathrm{d}t=\\ \int_{0}^{y_{i}}\lambda_{0}(t,\mathbf{x}_{i};\phi)\>\sigma(g(t,\mathbf{x}_{i};\boldsymbol{\theta}))\mathrm{d}t$ (50)
@@
Since $10<\sigma(\cdot)<$ 1, we have
@@
$\int_{0}^{y_{i}}\lambda_{0}(t,\mathbf{x}_{i};\phi)\sigma(g(t,\mathbf{x}_{i};\boldsymbol{\theta}))\mathrm{d}t<\infty.$ (51)
This shows the finiteness of the right-hand side of (50). By combining $\sigma(z)=1-\sigma(-z)$ with (23) we obtain that
$\int_{0}^{y_{i}}\lambda_{0}(t,\mathbf{x}_{i};\phi)\>\sigma(g(t,\mathbf{x}_{i};\boldsymbol{\theta}))\mathrm{d}t=\\ \int_{0}^{y_{i}}\int_{0}^{\infty}\left(1-e^{f(\omega,-g(t,\mathbf{x}_{i};\boldsymbol{\theta}))}\right)p_{\text{PG}}(\omega|1,0)\lambda_{0}(t,\mathbf{x}_{i};\phi)\mathrm{d}\omega\mathrm{d}t.$ (52)
Putting together the finiteness from (51) with the equality of (52) completes the proof. ∎
@@
By the Pólya–Gamma identity (Eq. (23)), the bracketed integral equals $\sigma(g(y_{i},\mathbf{x}_{i};\boldsymbol{\theta}))$. Multiplying by $\lambda_{0}(y_{i},\mathbf{x}_{i};\phi)^{\delta_{i}}$ gives exactly
@@
is absolutely convergent with probability 1, and by Lemma N.3 the corresponding integral converges. Therefore, we may apply Campbell’s Theorem (Theorem C.3) together with the PG-sigmoid identity from (50) to conclude

### v2

@@ Appendix I Computational Speed-Ups
Fix an arbitrary index $i\in\{1,\dots,N\}$. From Section 2.3, recall that $p_{\boldsymbol{\theta}}(\boldsymbol{\theta})$ is the probability density function of a multivariate normal distribution with zero mean and identity covariance matrix $\mathbf{I}_{m}$. Our goal is to show that the normalization factor $Z(t,\mathbf{x}_{i})$ admits a strictly positive lower bound on $0,y_{i}]$, from which the integrability of $\lambda_{0}(t,\mathbf{x}_{i};\phi)$ will follow.
Step 1: Continuity of $Z(t,\mathbf{x}_{i})$ on $0,y_{i}]$.
Fix any $t_{0}\in[0,y_{i}]$, and let $(t_{n})_{n\geq 1}$ be a sequence in $0,y_{i}]$ such that $t_{n}\to t_{0}$ as $n\to\infty$. Define, for each $n$, the functions
$h_{n}(\boldsymbol{\theta})$ $=\sigma(g(t_{n},\mathbf{x}_{i};\boldsymbol{\theta}))p_{\boldsymbol{\theta}}(\boldsymbol{\theta}),\quad n\geq 1,$
$h(\boldsymbol{\theta})$ $=\sigma(g(t_{0},\mathbf{x}_{i};\boldsymbol{\theta}))p_{\boldsymbol{\theta}}(\boldsymbol{\theta}).$
@@
since $0<\sigma(\cdot)<1$. Because $p_{\boldsymbol{\theta}}(\boldsymbol{\theta})$ integrates to 1 over $\mathbb{R}^{m}$, we may apply the Dominated Convergence Theorem (DCT) to conclude that:
@@
Since $t_{0}$ was arbitrary in $0,y_{i}]$, $Z$ is continuous everywhere on that interval.
Step 2: Strict positivity of $Z(t,\mathbf{x}_{i})$ on $0,y_{i}]$.
@@
Since $Z(t,\mathbf{x}_{i})$ is a continuous and strictly positive function on the compact interval $0,y_{i}]$, the Weierstrass Extreme Value Theorem ensures that $Z$ attains a minimum on this interval. Define:
@@
It is straightforward to verify that $\lambda_{0}(t;\phi)$ is integrable on $0,y_{i}]$, therefore it follows that
@@
$\mathbb{E}_{\omega\sim p_{\text{PG}}(\omega|1,0)}[\omega]=\frac{1}{4},$ (A32)
which follows by taking the limit $c\to 0$ in equation (A4). Alternatively, to prove (A32), one can start from the representation in equation (A1), apply Tonelli’s theorem to interchange expectation and infinite summation, and then invoke the series identity
@@
$\int_{0}^{y_{i}}\int_{0}^{\infty}\min(|f(\omega,-g(t,\mathbf{x}_{i};\boldsymbol{\theta}))|,1)\lambda_{i}(t,\omega;\phi)\mathrm{d}\omega\mathrm{d}t<\infty.$ (A33)
@@
$\mathcal{I}_{1}$ $=\int_{0}^{y_{i}}\int_{0}^{\infty}\frac{|g(t,\mathbf{x}_{i};\boldsymbol{\theta})|}{2}\lambda_{i}(t,\omega;\phi)\mathrm{d}t\mathrm{d}\omega,$
$\mathcal{I}_{2}$ $=\int_{0}^{y_{i}}\int_{0}^{\infty}\frac{g(t,\mathbf{x}_{i};\boldsymbol{\theta})^{2}}{2}\omega\lambda_{i}(t,\omega;\phi)\mathrm{d}\omega\mathrm{d}t,$
$\mathcal{I}_{3}$ $=\log(2)\int_{0}^{y_{i}}\int_{0}^{\infty}\lambda_{i}(t,\omega;\phi)\mathrm{d}\omega\mathrm{d}t.$
@@
Since $g(t,\mathbf{x}_{i};\boldsymbol{\theta})$ is continuous on the compact interval $0,y_{i}]$, it is bounded by some $M>0$. Then,
@@
Likewise $g(t,\mathbf{x}_{i};\boldsymbol{\theta})^{2}$ is bounded by some $C>0$ over $0,y_{i}]$ and $\mathbb{E}_{\omega\sim p_{\text{PG}}(\omega|1,0)}[\omega]=\frac{1}{4}$ (see  (A32)), so
@@
Since $\mathcal{I}_{1},\mathcal{I}_{2},\mathcal{I}_{3}$, are all finite, the condition in (A33) is satisfied and the sum $H(\Psi_{i})$ converges absolutely with probability 1. ∎ The next result presents an integral identity which is key to proving the data augmentation scheme of Theorem 3.1.
@@
$\int_{0}^{y_{i}}\int_{0}^{\infty}\left(1-e^{f(\omega,-g(t,\mathbf{x}_{i};\boldsymbol{\theta}))}\right)p_{\text{PG}}(\omega|1,0)\lambda_{0}(t,\mathbf{x}_{i};\phi)\mathrm{d}\omega\mathrm{d}t=\\ \int_{0}^{y_{i}}\lambda_{0}(t,\mathbf{x}_{i};\phi)\>\sigma(g(t,\mathbf{x}_{i};\boldsymbol{\theta}))\mathrm{d}t$ (A34)
@@
Since $0<\sigma(\cdot)<1$, we have
@@
$\int_{0}^{y_{i}}\lambda_{0}(t,\mathbf{x}_{i};\phi)\sigma(g(t,\mathbf{x}_{i};\boldsymbol{\theta}))\mathrm{d}t<\infty.$ (A35)
This shows the finiteness of the right-hand side of (A34). By combining $\sigma(z)=1-\sigma(-z)$ with (A5) we obtain that
$\int_{0}^{y_{i}}\lambda_{0}(t,\mathbf{x}_{i};\phi)\>\sigma(g(t,\mathbf{x}_{i};\boldsymbol{\theta}))\mathrm{d}t=\\ \int_{0}^{y_{i}}\int_{0}^{\infty}\left(1-e^{f(\omega,-g(t,\mathbf{x}_{i};\boldsymbol{\theta}))}\right)p_{\text{PG}}(\omega|1,0)\lambda_{0}(t,\mathbf{x}_{i};\phi)\mathrm{d}\omega\mathrm{d}t.$ (A36)
Putting together the finiteness from (A35) with the equality of (A36) completes the proof. ∎
@@
By the Pólya–Gamma identity (Eq. (A5)), the bracketed integral equals $\sigma(g(y_{i},\mathbf{x}_{i};\boldsymbol{\theta}))$. Multiplying by $\lambda_{0}(y_{i},\mathbf{x}_{i};\phi)^{\delta_{i}}$ gives exactly
@@
is absolutely convergent with probability 1, and by Lemma N.3 the corresponding integral converges. Therefore, we may apply Campbell’s Theorem (Theorem C.3) together with the PG-sigmoid identity from (A34) to conclude